40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 14637     Entry time: Fri May 24 17:50:19 2019
Author: gautam 
Type: Update 
Category: IOO 
Subject: IFO recovery 

At ~4pm, the main volume pressure (CC1) was reported to be ~5e-5 torr. So I replaced the HR mirror in the MC REFL path with the usual 10% beamsplitter, and aligned the beam onto MCREFL photodiode. I also replaced the ND filter on the AS port camera, and in front of the IPPOS QPD.

Then I turned up the power by HWP rotation - at the input to the IMC, I now measured 960 mW with the Coherent power meter, so the NPRO power has certainly decayed by ~10% from 2018 July. Normal high-power IMC autolocker script was re-enabled on megatron (and the slow servo enable threshold raised from 1000 cts to 8000cts). IMC was readily locked, after some hand alignment, I got a maximum of 14500 cts transmission. I was then able to lock the Y-arm. The dither alignment servo did not work with the nominal settings, but by hand alignment, I was able to get TRY up to 0.6 (I didn't try too hard to optimize this in any systematic way). X arm was also locked.

AUX drypump valved off and shutdown at ~610pm. I also switched both TP2 and TP3 to their lower rotation "standby" mode. So overall no major mishaps this time around. I am leaving the PSL shutter open over the long weekend. For in-air vs vacuum suspension spectra comparison, I kicked the ETMY optic at Fri May 24 18:26:10 PDT 2019.

ELOG V3.1.3-